# Efficient Agent Training for Computer Use

### v1

@@ Abstract
Scaling up high-quality trajectory data has long been a critical bottleneck for developing human-like computer use agents. We introduce PC Agent-E, an efficient agent training framework that significantly reduces reliance on large-scale human demonstrations. Starting with just 312 human-annotated computer use trajectories, we further improved data quality by synthesizing diverse action decisions with Claude 3.7 Sonnet. Trained on these enriched trajectories, our PC Agent-E model achieved a remarkable 141% relative improvement, surpassing the strong Claude 3.7 Sonnet with extended thinking on WindowsAgentArena-V2, an improved benchmark we also released. Furthermore, PC Agent-E demonstrates strong generalizability to different operating systems on OSWorld. Our findings suggest that strong computer use capabilities can be stimulated from a small amount of high-quality trajectory data. We open-source our entire suite of code, data, and models to facilitate future research at https://github.com/GAIR-NLP/PC-Agent-E.
11footnotetext: Equal Contribution.22footnotetext: Corresponding author.
@@ 1 Introduction
Developing autonomous agents that can operate computers as humans do Anthropic (2024); He et al. (2024); OpenAI (2025a); Qin et al. (2025); Team (2025) has long been a landmark pursuit in Artificial Intelligence (AI). Such agents could unlock unprecedented capabilities and significantly reduce human workload across various tasks. However, current models still fall significantly short of human performance Xie et al. (2024); Bonatti et al. (2024), largely due to insufficient computer knowledge and poor long-horizon planning ability. A key factor contributing to these deficiencies is the extreme scarcity of high-quality computer use trajectory data He et al. (2024); Ou et al. (2024); Xu et al. (2025a).
Previous works have investigated various strategies for trajectory collection. Some approaches rely entirely on human annotations to obtain trajectories Lu et al. (2024); Li et al. (2024), while others deploy agents to perform tasks directly in real-world environments Murty et al. (2024); Patel et al. (2024). In addition, some methods synthesize trajectories by leveraging data extracted from online tutorials Ou et al. (2024); Xu et al. (2025a).
@@
In this work, we explore efficient agent training for computer use. Inspired by recent findings Huang et al. (2024); Muennighoff et al. (2025); Ye et al. (2025) that synthesizing high-quality data using advanced reasoning models like Deepseek-R1 Guo et al. (2025) can efficiently enhance LLM reasoning, we extend the similar idea to the field of computer use agents.
@@
We propose PC Agent-E, an efficient agent training framework that integrates human expertise with AI automation. Starting from a small set of real-world human computer use trajectories, we leverage a frontier agent model to diversify action decisions (action with thought in ReAct Yao et al. (2023) paradigm), further enhancing the data quality. Training on these augmented trajectories, our agent demonstrates strong computer use capabilities with remarkable data efficiency.
@@
We begin by collecting 312 human computer use trajectories with PC Tracker He et al. (2024), a tool for gathering human-computer interaction data, with only two humans annotating one day. These trajectories include task descriptions, screenshots, and keyboard/mouse actions. We subsequently reconstruct the implicit thought process behind human actions, obtaining comprehensive human trajectories with thoughts. The successful completion of these tasks is inherently assured by human proficiency, obviating the need for additional verification.
@@
While these human trajectories already serve as valuable agent training data, we further enhance their quality through Trajectory Boost, a data synthesis method that enriches each trajectory step with a diverse set of alternative action decisions. The key insight is that computer use tasks can be completed through multiple valid pathways, meaning each step has various reasonable action alternatives supported by rational thought. To capture this diversity, we use a strong agent model to synthesize other possible action decisions. Specifically, recognizing that each human trajectory step captures an environment snapshot essential for computer use agents to make decisions, we provide these snapshots to Claude 3.7 Sonnet Anthropic (2025a) and sample multiple possible action decisions, thereby significantly enriching and diversifying the trajectory data.
Experimental results demonstrate that a small set of high-quality trajectory data can indeed elicit complex agent capabilities for computer use. Trained with only 312 trajectories augmented by Claude 3.7 Sonnet (without extended thinking), our PC Agent-E model achieves an impressive 141% relative improvement over the base model Qwen2.5-VL-72B and even outperforms Claude 3.7 Sonnet with extended thinking Anthropic (2025b) on WindowsAgentArena-V2, a benchmark we improved from WindowsAgentArena Bonatti et al. (2024). Furthermore, PC Agent-E demonstrates strong generalizability to different operating systems on OSWorld Xie et al. (2024), another popular computer use benchmark focused on Linux tasks. This extreme data efficiency directly stems from our focus on high-quality trajectories, ensuring both real-world task completion and fostering diverse action decisions through our human trajectory collection and Trajectory Boost method.
@@
Figure 2: Overview of our framework, consisting of four key components: (1) Trajectory Collection, gathering a small set of human trajectories by recording user actions and state observations at each step; (2) Thought Completion, reconstructing the implicit thought process missing in raw human trajectories; and (3) Trajectory Boost, diversifying action decisions to further enhance trajectory quality (4) Agent Training, developing a strong computer use agent with remarkable data efficiency.
@@
    We propose Trajectory Boost, a simple but effective data synthesis method that significantly enhances trajectory quality by enriching each step with diverse alternative action decisions. Leveraging each step of a human trajectory as an environment snapshot, our method substantially increases action decision diversity.
@@
    We release WindowsAgentArena-V2, an improved benchmark rectifying evaluation dependence, infeasible hacking and other limitations in the original WindowsAgentArena, ensuring more robust and fair evaluations of Windows computer use.
@@
    We developed PC Agent-E, a state-of-the-art (SOTA) open-source model for Windows computer use, which exemplifies exceptional data efficiency. Trained with just 312 augmented trajectories, our model successfully outperforms the strong Claude 3.7 Sonnet with extended thinking. Besides, we demonstrate its generalizability across different operating systems on OSWorld.
@@ 2.1 Computer Use Agent
With the advancement of Vision-Language Models (VLMs) Bai et al. (2025); Deitke et al. (2024), the way computer use agents interact with computers has gradually shifted from relying on textual representations such as accessibility trees Agashe et al. (2024); Wu et al. (2024a); Jia et al. (2024) to directly using screenshots Qin et al. (2025); Xu et al. (2025b); He et al. (2024). Existing computer use agents can be categorized based on the extent of human prior embedded in scaffold design: One is modular agent workflows Agashe et al. (2024); Wu et al. (2024a); Liu et al. (2025), which defines specialized modules and prompts multi-agents to collaborate. The other is native agent models Anthropic (2024); Qin et al. (2025); OpenAI (2025a), which depends on a single model to take action step by step based on its history and current state.
While modular agent workflows can reduce task complexity, their heavy reliance on human priors makes them vulnerable to unexpected errors and unfamiliar domains Qin et al. (2025). Furthermore, predefined workflows are hard to improve through end-to-end training Saltzer et al. (1984); Pan et al. (2024). With the continuous enhancement of model capabilities, native agent models have emerged as the dominant paradigm. This approach demonstrates superior flexibility and generalizability and offers sustainable improvement potential through either supervised fine-tuning (SFT) Xu et al. (2025b) or reinforcement learning (RL) OpenAI (2025a). Our work explores the efficient agent training methods for native agent models through SFT.
@@ 2.2 Data Synthesis
As Large Language Models (LLMs) grow ever more powerful OpenAI (2024a, b), it has become a common practice to use them to synthesize data. Distillation methods Taori et al. (2023); Gunasekar et al. (2023); Xu et al. (2023) leverage SOTA models to generate large-scale data for training smaller models. On the other hand, self-improvement methods enable a model to bootstrap and refine its own training data Wang et al. (2023). Beyond basic instruction-following,  Zelikman et al. (2022) first investigates synthesis thinking processes for reasoning tasks. More recently,  Muennighoff et al. (2025); Ye et al. (2025) demonstrate that advanced reasoning models can synthesize high-quality reasoning data that can boost a model’s reasoning capability efficiently.
In the domain of computer use agents, data synthesis can be broadly categorized into three aspects: (1) large-scale datasets that build foundational graphical user interface (GUI) understanding, with tasks like screenshot captioning or question–answering Liu et al. (2024); Qin et al. (2025) (2) single-step visual grounding tasks, where tasks of a single click are synthesized from specific locations on the GUI Gou et al. (2025); Wu et al. (2024b) (3) multi-step trajectory synthesis, in which recent research has explored leveraging web tutorials to guide trajectory generation Ou et al. (2024); Xu et al. (2025a) or reverse-synthesizes tasks from the agent’s own exploration records Sun et al. (2025); Murty et al. (2024). Our work differs from prior work by synthesizing high-quality multi-step trajectories from real-world human demonstrations and emphasizing data efficiency.
@@ 3.1 Overview
We propose PC Agent-E, an efficient agent training framework that elicits strong computer use capabilities with remarkable data efficiency, as illustrated in Figure 2. Our method generates high-quality trajectory data by combining authentic human-computer interactions with diverse action decisions, offering advantages in both realism and diversity.
@@ 3.2 Trajectory Collection
We collected human computer use trajectories (see an example at Figure 3) with PC Tracker He et al. (2024), a tool that records the screen state observation and the keyboard/mouse action at each step for a given task. The recorded actions are structured in a unified action space $\mathcal{A}$, as shown in Table 1. For task generation, we first manually compose a small seed set across multiple software applications and then enlarge it with LLMs. The resulting tasks were distributed to human annotators, who completed the tasks on their own Windows computers with PC Tracker recording trajectories automatically. After finishing a task, annotators could either discard unsatisfactory trajectories or modify the task descriptions based on their actual execution, thereby ensuring the correctness and completeness of the collected trajectories. We then applied a set of rule-based filters to remove entire trajectories or individual steps that exhibited errors or other undesirable behaviors.
We employed a rigorous data decontamination procedure on these collected trajectories. Each task description was compared against the tasks in our main evaluation benchmark (§4) using n-gram overlap and semantic similarity metrics. Trajectories with tasks exhibiting excessive similarity to any test task were removed from the dataset.
This procedure finally yielded 312 real-world human trajectories, with distribution across applications shown in Figure 4. The whole annotation process was completed by two annotators within a single day, with an average of roughly 3 minutes per trajectory. Given humans’ proficiency in computer use, the mechanisms for annotators to discard trajectories or revise task descriptions after execution, and our data filtering process, no additional verification was required to ensure trajectory correctness.
@@ 3.3 Thought Completion
We first reconstruct the implicit thought process behind human actions using an iterative approach. Specifically, for each action in the trajectory, we provide Claude 3.7 Sonnet with: task description, historical actions with their previously reconstructed thought processes, the current action, and the corresponding screenshot. Based on this information, the model generates human’s thought process behind the action. As shown in Figure 5, the recorded raw human trajectory was converted to a human trajectory with thoughts, where the reconstructed thought process is added to each step. See prompt in Appendix B.1.
@@ 3.4 Trajectory Boost
After thought completion, we obtain comprehensive human trajectories with explicit thought processes. While these trajectories already serve as valuable agent training samples, we further enhance their diversity through a simple but effective approach called Trajectory Boost, which synthesizes diverse alternative action decisions for each step of the trajectory.
The key insight behind Trajectory Boost is that computer use tasks inherently allow for multiple valid solution pathways. Consequently, at any given step, several reasonable actions supported by rational thought processes may exist, extending beyond the single solution adopted by human annotators. To capture this inherent diversity, we employ a frontier computer use agent, Claude 3.7 Sonnet, to generate these alternative action decisions. Its long-horizon planning capabilities, advanced reasoning patterns, and broad knowledge of computer use enable it to generate thought processes and actions that are highly informative and valuable, thereby substantially enhancing the richness and diversity of our trajectory data.
Specifically, we recognize that each step in a human trajectory captures an environment snapshot of the computer, providing the necessary information for both humans and agents to make decisions. For step $k$ on a human trajectory with observation $o_{k}$, thought process $t_{k}$, action $a_{k}$ and task description $T$, the environment snapshot is $<T,o_{k},h_{k}>$, where the history context $h_{k}=(t_{1},a_{1},t_{2},a_{2},\dots,t_{k-1},a_{t-1})$ is constructed with previous human steps. We input this environment snapshot to Claude 3.7 Sonnet instantiate in the PC Agent-E scaffold (§3.5), and sample multiple single-step action decisions $(t_{k}^{\prime}$, $a_{k}^{\prime})$ from it. Prompt used are shown in Appendix B.2. In practice, we sample 9 action decisions in parallel. This produces a Traj Tree, as shown in Figure 5, with human trajectory forming the main trunk and the augmented action decisions branching off as leaf nodes. These sampled action decisions from Claude 3.7 Sonnet are not executed in real computer environments, but serve as important augmented data for later agent training.
@@ 3.5 Agent Training
PC Agent-E adopts a deliberately simple end-to-end scaffold, as our primary focus is on validating the efficiency of our agent training methodology rather than optimizing performance through complex workflow design or elaborate prompt engineering. At inference, PC Agent-E takes a ⟨screenshot, task description, history⟩ input and outputs a ⟨thought, action⟩ decision in the ReAct Yao et al. (2023) paradigm, as shown in Figure 6. The action space is the same as $\mathcal{A}$ in Table 1, and every action is executed via the PyAutoGUI library. The history is a textual log of previous thoughts and actions. To maintain simplicity in both training and inference, past screenshots are excluded, although we believe that adding this image history would be beneficial for improving model performance. Prompt used for the scaffold is shown in Appendix B.3.
For training, we transform each action node from our Traj Tree into an individual training instance. The training sample structure and the inference-time scaffold of the agent share a direct correspondence, as illustrated in Figure 6. For both human-demonstrated and model-synthesized action nodes, the history in the training sample includes only prior human actions on the main trunk of the Trajectory Tree. This is consistent with the historical context available to both humans and the model when making the corresponding action decision. We finally obtained 27K training instances from 312 augmented trajectories, each following a consistent input-output structure at inference time.

### v2

@@ Abstract
Scaling up high-quality trajectory data has long been a critical bottleneck for developing human-like computer use agents. We introduce PC Agent-E, an efficient agent training framework that significantly reduces reliance on large-scale human demonstrations. Starting with just 312 human-annotated computer use trajectories, we further augment them by synthesizing diverse alternative action decisions with Claude 3.7 Sonnet. Trained on these enriched trajectories, our PC Agent-E model achieved a remarkable 141% relative improvement, and even surpassed the Claude 3.7 Sonnet by 10% in relative terms on WindowsAgentArena-V2, an improved benchmark we also released. By integrating robust human computer use skills with automated AI data synthesis capabilities, our method not only brought substantial improvements over training on human trajectories alone, but also significantly surpassed direct distillation from Claude 3.7 Sonnet. Code, data and models are available at https://github.com/GAIR-NLP/PC-Agent-E.
@@ 1 Introduction
Developing autonomous agents that can operate computers as humans do Anthropic (2024); OpenAI (2025) has long been a landmark pursuit in Artificial Intelligence (AI). Such computer use agents, powered by Vision-Language Models (VLM), perceive screenshots to interact directly with Graphical User Interfaces (GUI) — clicking buttons, navigating menus, and entering text. This allows them to automate a wide range of digital tasks, ranging from routine paperwork and online shopping to complex content creation, promising a significant reduction in manual human workload.
However, current models still fall significantly short of human performance Xie et al. (2024); Bonatti et al. (2024). This capability gap is even more pronounced within the open-source community, which lacks any solution competitive with leading proprietary systems like Claude 3.7 Sonnet Anthropic (2025a). Instilling these advanced computer use capabilities in open-source models remains an unsolved problem. A key factor contributing to these deficiencies is the extreme scarcity of high-quality computer use trajectory data Ou et al. (2024); Xu et al. (2025a).
@@
In this work, we explore efficient agent training for computer use, enabling open-source models to even exceed the performance of proprietary counterparts with minimal human annotation. Inspired by recent findings Huang et al. (2024); Muennighoff et al. (2025); Ye et al. (2025) that synthesizing high-quality data using advanced reasoning models like Deepseek-R1 Guo et al. (2025) can efficiently enhance LLM reasoning, we extend the similar idea to the field of computer use agents.
@@
We propose PC Agent-E, an efficient agent training framework that integrates human expertise with AI automation. Starting from a small set of real-world human computer use trajectories, we leverage a frontier agent model to diversify action decisions, yielding richer supervision signals. Training on these augmented trajectories, our agent demonstrates strong computer use capabilities with remarkable data efficiency.
@@
We begin by collecting 312 human computer use trajectories with PC Tracker He et al. (2024), a tool for gathering human-computer interaction data, with only two humans annotating one day. These trajectories include task descriptions, screenshots, and human keyboard/mouse actions. We subsequently reconstruct the implicit thought process behind human actions, obtaining comprehensive human trajectories with thoughts. The successful completion of these tasks is inherently assured by human proficiency, obviating the need for additional verification.
@@
While these human trajectories already serve as valuable agent training data, we further augment them through Trajectory Boost, a data synthesis method we developed to enrich each trajectory step with diverse alternative action decisions. The key insight is that computer use tasks can be completed through multiple valid pathways, meaning each step has various reasonable action alternatives supported by rational thought. To capture this diversity, we use a strong agent model to synthesize other possible action decisions. Specifically, each human trajectory step provides an environment snapshot that encapsulates the essential state information necessary for computer use agents to make decisions. We provide these snapshots to Claude 3.7 Sonnet and sample multiple possible action decisions, thereby significantly enriching and diversifying the trajectory data.
Experimental results demonstrate that with only a small set of human-annotated trajectories, our method can boost the performance of an open-source model to that of frontier models. Trained with only 312 trajectories augmented by Claude 3.7 Sonnet, our PC Agent-E model achieves an impressive 141% relative improvement over the base model Qwen2.5-VL-72B and even outperforms the teacher model Claude 3.7 Sonnet by 10% in relative terms on WindowsAgentArena-V2, a benchmark we improved from WindowsAgentArena Bonatti et al. (2024). Furthermore, PC Agent-E generalizes well to different operating systems on OSWorld Xie et al. (2024). Our ablation study further demonstrates that our method for utilizing human demonstrations is not only superior to relying solely on the human trajectories but is also significantly more effective and efficient than directly distilling from the teacher model.
@@
Figure 2: Overview of our framework, consisting of four key components: (1) Trajectory Collection, gathering a small set of human trajectories by recording user actions and state observations at each step; (2) Thought Completion, reconstructing the implicit thought process missing in raw human trajectories; and (3) Trajectory Boost, diversifying action decisions to further augment trajectories (4) Agent Training, developing a strong computer use agent with remarkable data efficiency.
@@
    We propose Trajectory Boost, a simple data synthesis method that unlocks remarkable data efficiency for training computer use agents. By augmenting human trajectories with diverse action decisions from a frontier model, our method demonstrates significantly greater effectiveness and efficiency than both using human data alone and direct distillation.
@@
    We release WindowsAgentArena-V2, an improved benchmark rectifying evaluation dependence, infeasible hacking, and other limitations in the original WindowsAgentArena benchmark, ensuring more robust and fair evaluations of computer use.
@@
    We developed PC Agent-E, an open-source computer use agent that achieves performance comparable to leading proprietary models. Trained with just 312 augmented trajectories, our model successfully outperforms the strong teacher model, Claude 3.7 Sonnet, showing exceptional data efficiency.
@@ 2.1 Computer Use Agent
With the advancement of VLMs Bai et al. (2025); Deitke et al. (2024), the way computer use agents interact with computers has gradually shifted from relying on textual representations such as accessibility trees Agashe et al. (2024); Wu et al. (2024a) to directly using screenshots Qin et al. (2025); Xu et al. (2025b); He et al. (2024). Existing computer use agents can be categorized based on how much human prior is built into their design: One is modular agent workflows Agashe et al. (2024); Wu et al. (2024a), which defines specialized modules and prompts multi-agents to collaborate. The other is native agent models Anthropic (2024); Qin et al. (2025); OpenAI (2025), which depends on a single model to take action step by step based on its history and current state.
While modular agent workflows can reduce task complexity, their heavy reliance on human priors hinders adaptation to new domains and limits end-to-end optimization Saltzer et al. (1984); Pan et al. (2024). With the continuous enhancement of model capabilities, native agent models have emerged as the dominant paradigm. This approach offers flexibility, generalizability, and sustainable gains via supervised fine-tuning (SFT) Xu et al. (2025b) or reinforcement learning (RL) OpenAI (2025). Our work explores the efficient agent training methods for native agent models through SFT.
@@ 2.2 Data Synthesis
As Large Language Models (LLMs) grow ever more powerful, it has become a common practice to use them to synthesize data. Distillation methods Taori et al. (2023); Gunasekar et al. (2023); Xu et al. (2023) leverage state-of-the-art (SOTA) models to generate large-scale data for training weaker models. On the other hand, self-improvement methods enable a model to bootstrap and refine its own training data Wang et al. (2023).
In the domain of computer use agents, data synthesis can be broadly categorized into three aspects: (1) large-scale datasets that build foundational GUI understanding, with tasks like screenshot captioning or question–answering Liu et al. (2024); Qin et al. (2025) (2) single-step visual grounding, where mouse click tasks are synthesized from specific locations on the GUI Gou et al. (2025); Wu et al. (2024b) (3) multi-step trajectory, in which recent research has explored leveraging web tutorials to guide trajectory generation Ou et al. (2024); Xu et al. (2025a) or reverse-synthesizes tasks from the agents’ own exploration records Sun et al. (2025); Murty et al. (2024). Our work differs from prior work by synthesizing high-quality multi-step trajectories based on real-world human demonstrations and emphasizing data efficiency.
@@ 3.1 Overview
We propose PC Agent-E, an efficient agent training framework for computer use that integrates human expertise with AI automation, as illustrated in Figure 2. Our method generates high-quality trajectory data by combining authentic human-computer interactions with diverse action decisions, offering advantages in both realism and diversity.
@@ 3.2 Trajectory Collection
Length Range Count Percentage
2–5 74 23.7%
6–10 139 44.6%
11–15 73 23.4%
16–30 26 8.3%
Table 1: Distribution of the 312 task
trajectories across length ranges (steps).
We collected human computer use trajectories with PC Tracker He et al. (2024), a tool that records the screen state observation and the human keyboard/mouse action at each step for a given task. The recorded actions are structured in a unified action space $\mathcal{A}$, as shown in Table 2. For task generation, we first manually compose a small seed set across multiple software applications and then enlarge it with LLMs. The resulting tasks were distributed to human annotators, who completed the tasks on their own Windows computers with PC Tracker recording trajectories automatically. After finishing a task, annotators could either discard unsatisfactory trajectories or modify the task descriptions based on their actual execution, thereby ensuring the correctness and completeness of the collected trajectories. We then applied a set of rule-based filters to remove entire trajectories or individual steps that exhibited errors or other undesirable behaviors.
We employed a rigorous data decontamination procedure on these collected trajectories. Each task description was compared against the tasks in our main evaluation benchmark (§4) using n-gram overlap and semantic similarity metrics. Specifically, we retained only trajectories whose task descriptions have an overlap of 0 under 13-gram and less than 0.7 under 3-gram, and a cosine semantic similarity under 0.85 with respect to any test task.
This procedure finally yielded 312 real-world human computer use trajectories, with distribution across applications shown in Figure 3. We report the trajectory length distribution in Table 1. No explicit instructions on trajectory length were given to annotators; the distribution naturally emerges from the collected tasks and human demonstration behavior. The whole annotation process was completed by two annotators within a single day, with an average of roughly 3 minutes per trajectory. Given humans’ proficiency in computer use, the mechanisms for annotators to discard trajectories or revise task descriptions after execution, and our data filtering process, no additional verification was required to ensure trajectory correctness.
@@ 3.3 Thought Completion
We first reconstruct the implicit thought process behind human actions using an iterative approach. Specifically, for each action in the raw trajectory, we provide Claude 3.7 Sonnet with: task description, historical actions with their previously reconstructed thought processes, the current action, and the corresponding screenshot. Based on this information, the model generates the implicit thought process behind action. As shown in Figure 4, the recorded raw human trajectory was converted to a human trajectory with thoughts, where the reconstructed thought process is added to each step. See our prompt in Appendix F.1.
@@ 3.4 Trajectory Boost
After thought completion, we obtain comprehensive human trajectories with explicit thought processes. While these trajectories already serve as valuable agent training samples, we further augment them through a simple but effective approach called Trajectory Boost, which synthesizes diverse alternative action decisions for each step of the trajectory.
The motivation behind Trajectory Boost is that computer use tasks inherently allow for multiple valid solution pathways. Consequently, at any given step, several reasonable actions supported by rational thought processes may exist, extending beyond the single solution adopted by human annotators. To capture this inherent diversity, we utilize a frontier computer use agent model, Claude 3.7 Sonnet, to generate single-step alternative action decisions. Its long-horizon planning capabilities, advanced reasoning patterns, and broad knowledge of computer use enable it to generate thought processes and actions that are highly informative and valuable, thereby substantially enhancing the richness and diversity of our trajectory data.
Specifically, we recognize that each step in a human trajectory captures an environment snapshot of the computer, providing the necessary information for both humans and agents to make decisions. For step $k$ on a human trajectory with observation $o_{k}$, thought process $t_{k}$, action $a_{k}$ and task description $T$, the environment snapshot is $T,o_{k},h_{k}>$, where the history context $h_{k}=(t_{1},a_{1},t_{2},a_{2},\dots,t_{k-1},a_{t-1})$ is constructed with previous human steps. We input this environment snapshot to Claude 3.7 Sonnet instantiate in the PC Agent-E scaffold (§3.5), and sample multiple single-step action decisions $(t_{k}^{\prime}$, $a_{k}^{\prime})$ from it. Prompts used are shown in Appendix LABEL:appendix:_traj_boost. In practice, we sample 9 action decisions in parallel. This produces a Traj Tree, as shown in Figure 4, with human trajectory forming the main trunk and the augmented action decisions branching off as leaf nodes. These sampled action decisions from Claude 3.7 Sonnet are not executed in real computer environments, but serve as important augmented data for later agent training.
@@ 3.5 Agent Training
PC Agent-E adopts a deliberately simple end-to-end scaffold, as our primary focus is on validating the effectiveness of our agent training methodology rather than optimizing performance through complex workflow design or elaborate prompt engineering. At inference, PC Agent-E takes ⟨screenshot, task description, history⟩ as input and outputs a ⟨thought, action⟩ decision in the ReAct Yao et al. (2023) paradigm, as shown in Figure 5. The action space is the same as $\mathcal{A}$ in Table 2, and every action is executed via the PyAutoGUI library. The history is a textual log of previous thoughts and actions. To maintain simplicity in both training and inference, past screenshots are excluded, although we believe that adding this image history would be beneficial for improving model performance. The prompt used for the scaffold is shown in Appendix LABEL:appendix:_scaffold.
For training, we transform each action node from our Traj Tree into an individual training sample. The training sample structure and the inference-time scaffold of the agent share a direct correspondence, as illustrated in Figure 5. For both human-demonstrated and model-synthesized action nodes, the history in the training sample includes only prior human actions on the main trunk of the Traj Tree. This is consistent with the historical context available to both humans and the model when making the corresponding action decision. We finally obtained 27K training samples from 312 augmented trajectories, each following a consistent input-output structure at inference time.